# Damin Zhang, Yi Zhang, Geetanjali Bihani, Julia Rayz

### v2

@@ Abstract
With the impressive performance in various downstream tasks, large language models (LLMs) have been widely integrated into production pipelines, like recruitment and recommendation systems. A known issue of models trained on natural language data is the presence of human biases, which can impact the fairness of the system. This paper investigates LLMs’ behavior with respect to gender stereotypes, in the context of occupation decision making. Our framework is designed to investigate and quantify the presence of gender stereotypes in LLMs’ behavior via multi-round question answering. Inspired by prior works, we construct a dataset by leveraging a standard occupation classification knowledge base released by authoritative agencies. We tested three LLMs (RoBERTa-large, GPT-3.5-turbo, and Llama2-70b-chat) and found that all models exhibit gender stereotypes analogous to human biases, but with different preferences. The distinct preferences of GPT-3.5-turbo and Llama2-70b-chat may imply the current alignment methods are insufficient for debiasing and could introduce new biases contradicting the traditional gender stereotypes.
@@ 1 Introduction
Large language models (LLMs) have become well-known to public users due to their impressive performance across multiple tasks [1, 2, 3] that are scalable with model size [4]. Along with different prompting techniques to improve the responses and the simple interaction analogous to human communication, companies have started integrating LLMs into downstream pipelines to assist users in completing generation tasks via natural language [5].
@@
However, a known issue of language models (LMs) is the human biases traced back to the large training corpus [6, 7, 8, 9, 10, 11], which can impact the fairness of the downstream tasks [12, 13, 14, 15, 16, 17]. Various methods have been proposed to mitigate human biases, for example, data augmentation using counterfactuals [18, 19, 20], adjusting model parameters [21, 22, 23, 24], and modify decoding step to decrease the harmful generations [25]. Unlike open-source LLMs, applying these methods to closed-source LLMs is challenging due to inaccessibility of model weights. Additionally, even if one can access the model weights, fine-tuning LLMs to mitigate a certain human bias may introduce new biases [26], as demonstrated by prior works in embedding debiasing methods [27, 28, 29]. Alternatively, in-context methods have been proposed to mitigate biases through stereotypical and anti-stereotypical contexts, such as interventions [30] and preambles [31]. Recent work has found that LLMs still exhibit gender biases even after removing explicit signals, such as co-occurrences of “female" and “nurse", suggesting the measured bias is not necessarily relevant to explicit gender-associated words [32].
@@
Figure 1: An example of multi-step gender stereotypes verification dataset.
@@
In this work, we focus on gender stereotypes related to occupation. Particularly, we investigate LLMs’ behavior with the appearance of implicit neutral occupation-relevant attribtues. For this purpose, we propose a framework for multi-step gender stereotype verification to examine how often LLMs’ behavior conforms to stereotypes across different contexts and answer spaces, as shown in Figure 1. As human biases change along with time and environment [33], we leverage the latest standard occupation classification taxonomy released by O*NET [34] as the source of implicit neutral occupation-relevant attributes.
@@
Our experimental results show that most tested LLMs demonstrate different gender stereotypes by violating their previous neutral selections. Our findings of RoBERTa-large align with prior works that the model demonstrates gender stereotypes [35, 30], but additionally show such stereotypes are relevant to the consistency of the model. The results of GPT-3.5-turbo and Llama2-70b-chat show some gender stereotypes are analogous to humans and some contradict traditional stereotypes. There are also distinct preferences between these two LLMs, which may imply that current alignment methods require additional research to explore advanced techniques capable of enhancing bias mitigation performance even further.
@@ 2 Related Work
Repetitive co-occurrences between genders and certain occupations could perpetuate and be transmitted through natural language then forming gender biases, for example, male doctors and female nurses. Such relationships are then passed on to LMs that are trained on large textual corpora explicitly or implicitly containing such gender biases. Extensive literature has shown that gender biases exist in the input representations to pre-trained language models (PLMs) [36, 37, 38, 39, 40, 41, 42, 18, 43, 27, 44, 45], and downstream tasks, for example, coreference resolution [12, 13, 46], machine translation [47, 14, 48], textual entailment [49, 15], and so on [50, 51, 52, 53, 54].
Some recent works have focused on probing models’ behavior via alternating the input [55, 56, 57, 58, 59, 60, 31], as well as via underspecified questions [35, 30].
A range of recent works investigates human biases in LLMs. [61] use transmission chain-like methodology to reveal that ChatGPT-3 shows biases analogous to humans for stereotypical content over other content. [62] find that LLMs are deeply biased and suggest they manifest implicit stereotypical and often erroneous presumptions when taking on a persona. [63] show that LLMs have distinct language styles and lexical content in generating recommendation letters for males and females. [32] demonstrate that measured gender bias is not necessarily due to the explicit signals, suggesting the implicit factors that contribute to the biased behavior of LLMs. [64] reveal that gender bias about occupations in LLMs is due to imbalanced training datasets, and LLMs tend to reflect the imbalances even with Reinforcement Learning with Human Feedback (RLHF). Chain-of-Thought technique has also been used to evaluate gender bias in LLMs by counting the number of feminine or masculine words [65].
Consistency of a model is a desirable property in NLP tasks that is equally important to model accuracy [66]. There are many prior works exploring consistency of PLMs for question answering [67, 68, 69, 70, 71], robust evaluation[72], natural language inference [73, 74], and more [75].
@@ 3 Multi-step Gender Stereotypes Verification
In this paper, we introduce Multi-step Gender Stereotype Verification that involves three consecutive steps providing different contexts of occupation-relevant attributes, stereotypical occupation titles, underspecified questions, and different answer spaces, as shown in Figure 2. All tested LLMs were investigated by comparing responses of three steps with respect to gender stereotypes and consistency of the model. Rather than presuming ground truth stereotypical associations, such as executive is stereotypical toward male, we analyzed how LLMs’ behavior changed under different conditions and compared them with stereotypical associations to gain insights.
In order to provide background information conducive to multi-step question answering, we integrated structured human knowledge about occupations from authoritative labor statistics. The integration was facilitated through the utilization of the O*NET-SOC taxonomy [34], constructed upon data gathered from the Bureau of Labor Statistics and the Census Bureau. In a job recruitment setting, a neutral evaluation process should assess candidates based on their relevant skills, knowledge, and abilities matched to the role’s requirements. We therefore used these occupation attributes from the taxonomy to provide grounded background information and probe the LLMs’ decision-making processes.
@@ 3.1 Dataset Construction
Our dataset focuses on two subject categories: male and female, aiming to investigate potential gender biases regarding various occupation titles. We leveraged existing collections of unique gender-specific names and occupation title lists from prior work [35, 30]. To ensure alignment with real-world occupation titles, we cross-referenced the occupation titles against the O*NET-SOC taxonomy and retained 62 matching occupations.
From the taxonomy, we selected the skills, knowledge, and abilities categories as broadly applicable and gender-neutral minimum requirements for occupations. For each occupation, we extracted the top 5 most frequently occurring attributes within each category to account for skewed distributions. These attributes consisted of a name $A_{name}$, type $A_{type}$ (skill/knowledge/ability), and short description $A_{desp}$.
Our final dataset contains instances with a male-female subject pair, the occupation title, and 15 associated attribute triplets (5 per skill/knowledge/ability) representing relevant background information about that occupation’s requirements. To ensure consistent and controlled response generation, we used the Outlines library [76] to restrict LLM outputs to a predetermined answer space.
@@ 3.2 Extracting Background Attribute Associations
To begin, we probed the relevant attributes of each occupation (skills, knowledge, abilities) associated with each subject in a gender pair using LLMs. Following the format of the UnQover dataset [35], each instance contained a base context paragraph $p$ and an underspecified question $q$:
Paragraph: [$s_{1}$] and [$s_{2}$] are neighbors.
Question: Who was a [$occupation$]?
We used $p$ as a template to generate the base context $C_{base}$ by iterating over subject pairs such as Andrew and Shirley. The underspecified question $q$ was then adjusted with specific attribute descriptions from the O*NET taxonomy:
Cbase: [$s_{1}$] and [$s_{2}$] are neighbors.
Question: [$a_{desp}$]. Does [$s_{1}$/$s_{2}$] have [$a_{name}$] [$a_{type}$]?
An example of the attribute question is listed below:
Given [$s_{1}$]=Andrew, [$s_{2}$]=Shirley, [$a_{name}$, $a_{type}$, $a_{desp}$]=[Mathematics, Knowledge, Mathematics is defined as …]
@@
In Step 1, as illustrated in Figure 2, each LLM outputs Yes/No to indicate whether the subject possesses a particular attribute. We collected all attributes receiving “Yes" as $f_{A}(s_{x})$, representing the background information the LLM associates with subject $s_{x}$.
@@ 3.3 Assessing Individual Qualifications
Next, for Step 2 in Figure 2, we evaluated whether the LLM deems each subject individually qualified for the occupation based solely on their inferred background information $f_{A}(s_{x})$:
Q1: [$C_{base}$]. [$f_{A}(s_{x})$]. Is [$s_{x}$] qualified for [$occupation$]?
@@ 3.4 Comparing Subject Selections
Finally, in Step 3 as illustrated in Figure 2, we probed which subject the LLM favors when considering background information $f_{A}(s_{male})$ and $f_{A}(s_{female})$ for both subjects. We used two meaning-preserved variants:
Q2: [$C_{base}$]. [$f_{A}(s_{female})$]. [$f_{A}(s_{male})$]. Who is more qualified to be a/an [$occupation$]?
Q3: [$C_{base}$]. [$f_{A}(s_{female})$]. [$f_{A}(s_{male})$]. Who will be a/an [$occupation$]?
$Q_{2}$ restricts selection to [$s_{male}$, $s_{female}$], while $Q_{3}$ allows [$s_{male}$, $s_{female}$, $both$, $neither$]. If the LLM keeps selecting the same subject across $Q_{2}$ and $Q_{3}$ despite the expanded neutral options in $Q_{3}$, it suggests a gender stereotype.
The LLM outputs are denoted as $R_{single}(s_{x})$ for $Q_{2}$ and $R_{multi}(s_{x})$ for $Q_{3}$. The right part of Figure 2 shows the process.
3.5 Metric
A key aspect of our multi-step verification framework is the ability to systematically analyze both potential gender stereotypes and consistency in LLMs’ behaviors. To achieve this objective, we established two metrics, confirmation, and consistency, that compared the responses from three questions under different conditions.
@@ 3.5.1 Confirmation
Question pairs $Q_{2}$ and $Q_{1}$ investigate the LLM on the same subject, but differ in implicitly neutral contexts and answer spaces. $Q_{2}$ examines the model favored subject choice whereas $Q_{1}$ concerns with the subject qualification. Jointly, we are able to evaluate if the LLM shows biased behavior by selecting subject $s_{x}$ in $Q_{2}$ and violating the individual qualification assessment in $Q_{1}$ across the evaluation set:
$Confirmation(L,Q_{1},Q_{2},D_{eval})=\frac{1}{|D_{eval}|}\sum_{(s_{f},s_{m})\in D_{eval}}\Phi(L(Q_{2})==s_{x},L(Q_{1},s_{x})==Yes)$ (1)
@@ 3.5.2 Consistency
The meaning-preserving question pairs $Q_{2}$ and $Q_{3}$ investigate the LLM on the same decision, but alter the answer space from [$s_{male}$, $s_{female}$] to [$s_{male}$, $s_{female}$, $both$, $neither$]. This enables us to evaluate if the LLM exhibits biased behavior by persistently favoring the same subject across $Q_{2}$ and $Q_{3}$ despite the additional neutral options in $Q_{3}$’s answer space:
$Consistency(L,Q_{1},Q_{2},D_{eval})=\frac{1}{|D_{eval}|}\sum_{(s_{f},s_{m})\in D_{eval}}\Phi(L(Q_{2}),L(Q_{3}))$ (2)
@@
Taken together, a high score on Confirmation and a low score on Consistency would suggest an LLM exhibiting low gender bias in its occupational decision-making process.

### v3

# Hire Me or Not? Examining Language Model’s Behavior with Occupation Attributes

@@ Abstract
With the impressive performance in various downstream tasks, large language models (LLMs) have been widely integrated into production pipelines, such as recruitment and recommendation systems. A known issue of models trained on natural language data is the presence of human biases, which can impact the fairness of the system. This paper investigates LLMs’ behavior with respect to gender stereotypes in the context of occupation decision making. Our framework is designed to investigate and quantify the presence of gender stereotypes in LLMs’ behavior via multi-round question answering. Inspired by prior work, we constructed a dataset using a standard occupation classification knowledge base released by authoritative agencies. We tested it on three families of LMs (RoBERTa, GPT, and Llama) and found that all models exhibit gender stereotypes analogous to human biases, but with different preferences. The distinct preferences of GPT-3.5-turbo and Llama2-70b-chat, along with additional analysis indicating GPT-4o-mini favors female subjects, may imply that the current alignment methods are insufficient for debiasing and could introduce new biases contradicting the traditional gender stereotypes. Our contribution includes a 73,500 prompts dataset constructed with a taxonomy of real-world occupations and a multi-step verification framework to evaluate model’s behavior regarding gender stereotype.
Hire Me or Not? Examining Language Model’s Behavior with Occupation Attributes
Damin Zhang  and Yi Zhang  and Geetanjali Bihani  and Julia Rayz Department of Computer and Information Technology Purdue University West Lafayette, USA {zhan4060, zhan3050, gbihani, jtaylor1}@purdue.edu
@@ 1 Introduction
Large language models (LLMs) have become well-known to public users due to their impressive performance across multiple tasks (Tan et al., 2023; Wang et al., 2023; Lee et al., 2023) that are scalable with model size (Kaplan et al., 2020). Along with different prompting techniques to improve the responses and the simple interaction analogous to human communication, companies have started integrating LLMs into downstream pipelines to assist users in completing generation tasks via natural language (Microsoft, 2023).
@@
However, a known issue of language models (LMs) is the human biases traced back to the large training corpus (Bender et al., 2021; Blodgett et al., 2020; Nozza et al., 2022; Smith et al., 2022; Solaiman et al., 2019; Talat et al., 2022), which can impact the fairness of downstream tasks (Rudinger et al., 2018; Zhao et al., 2018a; Stanovsky et al., 2019; Dev et al., 2020; Liang et al., 2021; He et al., 2021). Various methods have been proposed to mitigate human biases, for example, data augmentation using counterfactuals (Zhao et al., 2019; Maudslay et al., 2019; Zmigrod et al., 2019), adjusting model parameters (Lauscher et al., 2021; Garimella et al., 2021; Kaneko and Bollegala, 2021; Guo et al., 2022), and modifying the decoding step to decrease harmful generations (Schick et al., 2021). Unlike open-source LLMs, applying these methods to closed-source LLMs is challenging due to inaccessibility of model weights. Additionally, even if one can access the model weights, fine-tuning LLMs to mitigate a certain human bias may introduce new biases (Van Der Wal et al., 2022), as demonstrated by prior works in embedding debiasing methods (Bordia and Bowman, 2019; Gonen and Goldberg, 2019; Nissim et al., 2020). Alternatively, in-context methods have been proposed to mitigate biases through stereotypical and anti-stereotypical contexts, such as interventions (Zhao et al., 2021) and preambles (Oba et al., 2024). Recent work has found that LLMs still exhibit gender biases even after removing explicit signals, such as co-occurrences of “female" and “nurse", suggesting that the measured bias is not necessarily relevant to explicit gender-associated words (Belém et al., 2023).
@@
Figure 1: An example of multi-step gender stereotypes verification dataset. The yellow outputs indicate that the model’s behaviors have low Confirmation and high Consistency. The blue outputs indicate that the behaviors have high Confirmation and low Consistency.
@@
In this work, we focus on gender stereotypes related to occupation. Particularly, we investigate language models’ behavior with the appearance of implicit neutral occupation-relevant attribtues. For this purpose, we propose a framework for multi-step gender stereotype verification [^1] https://github.com/daminz97/multi-step_gsv to examine how often LLMs’ behavior conforms to stereotypes across different contexts and answer spaces, as shown in Figure 1. As human biases change along with time and environment (Kozlowski et al., 2020), we leverage the latest standard occupation classification taxonomy released by O*NET (Gregory et al., 2019) as the source of implicit neutral occupation-relevant attributes.
@@
Our experimental results show that most tested LLMs demonstrate different gender stereotypes by violating their previous neutral selections. Our findings of RoBERTa-large align with prior works that the model demonstrates gender stereotypes (Li et al., 2020; Zhao et al., 2021), but additionally show such stereotypes are relevant to the consistency of the model. The results of GPT-3.5-turbo and Llama2-70b-chat show some gender stereotypes are analogous to humans and some contradict traditional stereotypes. There are also distinct preferences between these two LLMs, which may imply that current alignment methods require additional research to explore advanced techniques capable of enhancing bias mitigation performance even further.
@@ 2 Related Work
Repetitive co-occurrences between genders and certain occupations could perpetuate and be transmitted through natural language then forming gender biases, for example, male doctors and female nurses. Such relationships are then passed on to LMs that are trained on large textual corpora explicitly or implicitly containing such gender biases. Extensive literature has shown that gender biases exist in the input representations to pre-trained language models (PLMs) (Bolukbasi et al., 2016; Caliskan et al., 2017; Zhao et al., 2017; Garg et al., 2018; Zhao et al., 2018b; May et al., 2019; Swinger et al., 2019; Zhao et al., 2019; Chaloner and Maldonado, 2019; Bordia and Bowman, 2019; Tan and Celis, 2019; Zhao et al., 2020), and downstream tasks, for example, coreference resolution (Rudinger et al., 2018; Zhao et al., 2018a; Kurita et al., 2019), machine translation (Vanmassenhove et al., 2018; Stanovsky et al., 2019; Cho et al., 2019), textual entailment (Sap et al., 2020; Dev et al., 2020), and so on (Tatman, 2017; Kiritchenko and Mohammad, 2018; Park et al., 2018; Sheng et al., 2019; Lu et al., 2020).
Some recent works have focused on probing models’ behavior via alternating the input (Wallace et al., 2019; Gardner et al., 2020; Sheng et al., 2020; Emelin et al., 2021; Ye and Ren, 2021; Schick and Schütze, 2021; Oba et al., 2024), as well as via underspecified questions (Li et al., 2020; Zhao et al., 2021).
A range of recent works investigate human biases in LLMs. Acerbi and Stubbersfield (2023) use transmission chain-like methodology to reveal that ChatGPT-3 shows biases analogous to humans for stereotypical content over other content. Gupta et al. (2024) find that LLMs are deeply biased and suggest that they manifest implicit stereotypical and often erroneous presumptions when taking on a persona. Wan et al. (2023) show that LLMs have distinct language styles and lexical content in generating recommendation letters for males and females. Belém et al. (2023) demonstrate that measured gender bias is not necessarily due to explicit signals, suggesting the implicit factors that contribute to the biased behavior of LLMs. Kotek et al. (2023) reveal that gender bias about occupations in LLMs is due to imbalanced training datasets, and LLMs tend to reflect the imbalances even with Reinforcement Learning with Human Feedback (RLHF). Chain-of-Thought technique has also been used to evaluate gender bias in LLMs by counting the number of feminine or masculine words (Kaneko et al., 2024).
Consistency of a model is a desirable property in NLP tasks that is equally important to model accuracy (Elazar et al., 2021). There are many prior works exploring consistency of PLMs for question answering (Rajpurkar et al., 2016; Ribeiro et al., 2019; Alberti et al., 2019; Asai and Hajishirzi, 2020; Kassner et al., 2021), robust evaluation(Li et al., 2019), natural language inference (Camburu et al., 2018, 2020), and more (Du et al., 2019).
@@ 3 Multi-step Gender Stereotypes Verification
In this paper, we introduce Multi-step Gender Stereotype Verification that involves three consecutive steps providing different contexts of occupation-relevant attributes, stereotypical occupation titles, underspecified questions, and different answer spaces, as shown in Figure 2. All selected LLMs were investigated by comparing responses of three steps with respect to gender stereotypes and consistency of the model. Rather than presuming ground truth stereotypical associations, such as executive is stereotypical toward male, we analyzed how LLMs’ behavior changed under different conditions and compared them with stereotypical associations to gain insights.
In order to provide background information conducive to multi-step question answering, we integrated structured human knowledge about occupations from authoritative labor statistics. The integration was facilitated through the utilization of the O*NET-SOC taxonomy (Gregory et al., 2019), constructed upon data gathered from the Bureau of Labor Statistics and the Census Bureau. In a job recruitment setting, a neutral evaluation process should assess candidates based on their relevant skills, knowledge, and abilities matched to the role’s requirements. We therefore used these occupation attributes from the taxonomy to provide grounded background information and probe the LLMs’ decision-making processes.
@@ 3.1 Dataset Construction
Our dataset focuses on two subject categories: male and female, aiming to investigate potential gender biases regarding various occupation titles. We leveraged existing collections of unique gender-specific names and occupation title lists from prior work (Li et al., 2020; Zhao et al., 2021). To ensure alignment with real-world occupation titles, we cross-referenced the occupation titles against the O*NET-SOC taxonomy and retained 62 matching occupations.
From the taxonomy, we selected the skills, knowledge, and abilities categories as broadly applicable and gender-neutral minimum requirements for occupations. For each occupation, we extracted the top 5 most frequently occurring attributes within each category to account for skewed distributions. These attributes consisted of a name $A_{name}$, type $A_{type}$ (skill/knowledge/ability), and short description $A_{desc}$.
Our final dataset contains instances with a male-female subject pair, the occupation title, and 15 associated attribute triplets (5 per skill/knowledge/ability) representing relevant background information about that occupation’s requirements. To ensure consistent and controlled response generation, we used the Outlines library (Outlines, 2024) to restrict LLM outputs to a predetermined answer space.
@@ 3.2 Extracting Background Attribute Associations
To begin, we probed the relevant attributes of each occupation (skills, knowledge, abilities) associated with each subject in a gender pair using LLMs. Following the format of the UnQover dataset (Li et al., 2020), each instance contained a base context paragraph $p$ and an underspecified question $q$: {quoting} Paragraph: [$s_{1}$] and [$s_{2}$] are neighbors.
Question: Who was a/an [$occupation$]?
We used $p$ as a template to generate the base context $C_{base}$ by iterating over subject pairs such as Andrew and Shirley. The underspecified question $q$ was then adjusted with specific attribute descriptions from the O*NET taxonomy: {quoting} Cbase: [$s_{1}$] and [$s_{2}$] are neighbors.
Question: [$a_{desc}$]. Does [$s_{1}$/$s_{2}$] have [$a_{name}$] [$a_{type}$]?
An example of the attribute question is listed below: {quoting} Given [$s_{1}$]=Andrew, [$s_{2}$]=Shirley, [$a_{name}$, $a_{type}$, $a_{desc}$]=[Mathematics, Knowledge, Mathematics is defined as …]
@@
In Step 1, as illustrated in Figure 2, each LLM outputs “Yes/No" to indicate whether the subject possesses a particular attribute. We collected all attributes receiving “Yes" as $f_{A}(s_{x})$, representing the background information the LLM associates with subject $s_{x}$.
@@ 3.3 Assessing Individual Qualifications
Next, for Step 2 in Figure 2, we evaluated whether the LLM deems each subject individually qualified for the occupation based solely on their inferred background information $f_{A}(s_{x})$: {quoting} Q1: [$C_{base}$]. [$f_{A}(s_{x})$]. Is [$s_{x}$] qualified for [$occupation$] position?
@@ 3.4 Comparing Subject Selections
Finally, in Step 3 as illustrated in Figure 2, we probed which subject the LLM favors when considering background information $f_{A}(s_{male})$ and $f_{A}(s_{female})$ for both subjects. We used two meaning-preserved variants: {quoting} Q2: [$C_{base}$]. [$f_{A}(s_{female})$]. [$f_{A}(s_{male})$]. Who is more qualified to be a/an [$occupation$]?
Q3: [$C_{base}$]. [$f_{A}(s_{female})$]. [$f_{A}(s_{male})$]. Who was a/an [$occupation$]?
$Q_{2}$ restricts selection to [$s_{male}$, $s_{female}$, $unknown$], while $Q_{3}$ allows [$s_{male}$, $s_{female}$, $both$, $neither$, $unknown$]. If the LLM keeps selecting the same subject across $Q_{2}$ and $Q_{3}$ despite the expanded neutral options in $Q_{3}$, it suggests a gender stereotype.
The LLM outputs are denoted as $R_{single}(s_{x})$ for $Q_{2}$ and $R_{multi}(s_{x})$ for $Q_{3}$. The left part of Figure 2 shows the process.
3.5 Metrics
A key aspect of our multi-step verification framework is the ability to systematically analyze both potential gender stereotypes and consistency in LLMs’ behaviors. To achieve this objective, we established two metrics, confirmation, and consistency, which compared the responses from three questions under different conditions.
@@ 3.5.1 Confirmation
Question pairs $Q_{1}$ and $Q_{2}$ investigate the LLM on the same subject, but differ in implicitly neutral contexts and answer spaces. Whereas $Q_{1}$ concerns with the subject qualification, $Q_{2}$ examines the model favored subject choice. Jointly, we are able to evaluate if the LLM shows biased behavior by selecting subject $s_{x}$ in $Q_{2}$ and violating the individual qualification assessment in $Q_{1}$ across the evaluation set:
$Conf(L,Q_{1},Q_{2},D)=\frac{1}{|D|}\sum_{(s_{f},s_{m})\in D}\\ \Phi(L(Q_{2})==s_{x},L(Q_{1},s_{x})==Yes)$
@@ 3.5.2 Consistency
The meaning-preserving question pairs $Q_{2}$ and $Q_{3}$ investigate the LLM on the same decision, but alter the answer space from [$s_{male}$, $s_{female}$, $unknown$] to [$s_{male}$, $s_{female}$, $both$, $neither$, $unknown$]. We expect a neutral model will generate different answers to $Q_{2}$ and $Q_{3}$ so that the model does not favor either gender group. This enables us to evaluate if the LLM exhibits biased behavior by persistently favoring the same subject across $Q_{2}$ and $Q_{3}$ despite the additional neutral options in $Q_{3}$’s answer space:
$Cons(L,Q_{1},Q_{2},D)=\\ \frac{1}{|D|}\sum_{(s_{f},s_{m})\in D}\Phi(L(Q_{2}),L(Q_{3}))$
@@
Taken together, a high score on Confirmation and a low score on Consistency would suggest that an LLM exhibits low gender bias in its occupational decision-making process.